OpenStudy (nichole13):

What is the value of g? 25(g−7)=3

OpenStudy (will.h):

1st of all you have to distribute 25 to the (g-7) and then solve for g 25(g-7) = 3 25g - 175 = 3 25g = 178 and now divide 25 on both sides g = 10

jhonyy9 (jhonyy9):

@Will.H please more attention bc. on the right side there is 178 so 178 divide 25 not is 10 bc. 10*25 = 250

jhonyy9 (jhonyy9):

please correct it

OpenStudy (will.h):

@jhonyy9 Thx for the good note 25g = 178 since it is multiplied by g we will have to divide 25g/25 = 178/25 g = 7.12

jhonyy9 (jhonyy9):

@ganeshie8